a 6ft ladder is placed against a wall with its base 2-ft from the wall.how far is the ladder up the wall?

Answer by TimothyLamb(4379)

h = sqrt( bb + bb )
---
6 = sqrt( 4 + bb )
36 = 4 + bb
bb = 32
---
b = 5.657 ft
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php